OpenStudy (bookworm14):

Please Help!! No Idea how to work this out!! :( medal and fan if possible!!

OpenStudy (bookworm14):

If an object is dropped from a height of 195 feet, the function h(t) =−16t^2+195 gives the height of the object after t seconds. When will the object hit the ground? a. 3.49 s c. 13.96 s b. 1.75 s d. 6.98 s

OpenStudy (bookworm14):

h(t)= 16t^2 + 195

OpenStudy (dan815):

when it hits the ground its height is 0

OpenStudy (dan815):

0=16t^2+195

OpenStudy (dan815):

solve for time

OpenStudy (bookworm14):

I am not sure how :/

OpenStudy (bookworm14):

do i have to divide or something?

OpenStudy (bookworm14):

@dan815 please? I literally have no idea how to solve for the time

OpenStudy (dan815):

yeah we want to isolate for t

OpenStudy (dan815):

both sides of the equations are equal we can the same thing to both sides and it will still be equal

OpenStudy (dan815):

we can do*

OpenStudy (dan815):

0-195=-16t^2+195 -195 -195=-16t^2 -195/-16 = -16/-16 * t^2 (-195/-16)=t^2 t=sqrt(195/16)

OpenStudy (bookworm14):

ok, I did that the only part i didnt make it to was t=sqrt(195/16) 3.49?

OpenStudy (dan815):

here is a good place to check your answers http://www.wolframalpha.com/input/?i=0%3D-16t^2%2B195

OpenStudy (dan815):

3.49 is right

OpenStudy (bookworm14):

thank you, i have one more, but using what you just taught me, i already solved it. would you mind checking it just realy quick?

OpenStudy (bookworm14):

nevermind, i just used the wolfram alpha xP

OpenStudy (dan815):

good :)